OpenStudy (anonymous):

Express each of the following in terms of a reference angle. Explain, please. 1a. sin 45º 1b. 135º 1c. sin 225º 1d. sin 315º 3a. cos 224.5º 3b. cos 658º 3c. 145.7º 3d. sin (-201º)

OpenStudy (dumbcow):

Let A be the angle, R be reference angle Quad 1: R =A Quad 2: R = 180-A Quad 3: R = A-180 Quad 4: R = 360 -A

OpenStudy (dumbcow):

if angle greater than 360, subtract 360 first if angle is neg, add 360 first

OpenStudy (anonymous):

i don't understand....

OpenStudy (dumbcow):

what do you mean? sin(135) = sin(45) because 135 degrees is in 2nd Quadrant, 180-135 = 45